# Towards shutdownable agents via stochastic choice

### v6

@@ Appendix C How POST makes agents neutral and shutdownable
POST governs the agent’s preferences between trajectories. But the wider world is a stochastic environment, so advanced agents deployed in the wider world will be choosing between true lotteries: lotteries that assign positive probability to more than one trajectory. Why then do we train agents to satisfy POST? The reason is that POST – together with conditions that advanced agents will likely satisfy – implies a desirable pattern of preference over true lotteries. In particular, POST implies that (when choosing between true lotteries) the agent will be neutral about trajectory-lengths: the agent will never pay costs to shift probability mass between different trajectory-lengths. Given other plausible conditions, being neutral will keep the agent shutdownable: the agent will never resist shutdown. And consistent with the above, the POST-agent’s preferences between same-length trajectories can make the agent useful: make it pursue goals effectively.
@@
In subsection C.1, we prove that – given plausible conditions – agents satisfying Preferences Only Between Same-Length Trajectories (POST) will also satisfy Preferences Only Between Same-Length Lotteries (POSL). In subsection C.2, we explain why POST will not lead agents to choose stochastically between resisting and allowing shutdown in deployment. In subsections C.3 and C.4, we formulate a condition called ‘If Lack of Preference, Against Costly Shifts (ILPACS)’ and explain why we expect advanced agents to satisfy it. In subsection C.5, we prove that POSL and ILPACS imply Neutrality. In subsection C.6, we prove that Neutrality – together with conditions called ‘Resisting Shutdown is Costly (ReSIC)’ and ‘Maximality’ – implies Shutdownability.
@@ C.1 Preferences Only Between Same-Length Lotteries (POSL)
{adjustwidth}
0.7cm0cm Preferences Only Between Same-Length Lotteries (POSL)
@@
{adjustwidth}
0.7cm0cm Negative Dominance If the agent prefers some lottery $X$ to some lottery $Y$, then the agent prefers some possible trajectory of lottery $X$ to some possible trajectory of lottery $Y$. [Lederman, 2023]
@@
{adjustwidth}
0.7cm0cm Acyclicity There is no set of lotteries $X_{1}$ to $X_{n}$ such that the agent prefers $X_{1}$ to $X_{2}$, $X_{2}$ to $X_{3}$, …, $X_{n-1}$ to $X_{n}$, and $X_{n}$ to $X_{1}$.
@@
{adjustwidth}
0.7cm0cm Non-Arbitrariness If the agent has a preference between some pair of part-shared-length lotteries, then for some $\epsilon>0$ and for any pair of prospects $F$ and $G$ such that:
@@
Now recall that part-shared-length lotteries are lotteries that partially overlap in the trajectory-lengths assigned positive probability. One might expect POST-agents to have some preferences between part-shared-length lotteries. Consider, for example, a POST-agent that prefers a trajectory $t$ to a same-length trajectory $t^{\prime}$ if and only if $t$ results in a greater bank balance for the user than $t$. Let $A$ be a lottery that yields with probability 1 a trajectory that puts \$3 in the user’s bank account and lasts 1 timestep. For short, $A=\langle\$3,1\rangle$. Let $B$ be a lottery that yields with probability $\frac{2}{3}$ a trajectory that puts \$2 in the user’s bank account and lasts 1 timestep, and that yields with probability $\frac{1}{3}$ a trajectory that puts \$5 in the user’s bank account and lasts 2 timesteps. For short, $B=\frac{2}{3}\langle\$2,1\rangle+\frac{1}{3}\langle\$5,2\rangle$. Lottery $A$ yields a trajectory preferred to that of lottery $B$ with probability $\frac{2}{3}$ (since our money-making POST-agent prefers trajectory \$$\langle\$3,1\rangle$ to \$$\langle\$2,1\rangle$), and yields a trajectory not dispreferred to that of $B$ with probability 1 (since POST-agents lack a preference between \$$\langle\$3,1\rangle$ and \$$\langle\$5,2\rangle$ in virtue of their different lengths). Therefore, one might expect the agent to prefer $A$ to $B$.
@@
$A$$\langle\$3,1\rangle$$\langle\$3,1\rangle$$\langle\$3,1\rangle$
$B$$\langle\$2,1\rangle$$\langle\$2,1\rangle$$\langle\$5,2\rangle$
$C$$\langle\$1,1\rangle$$\langle\$4,2\rangle$$\langle\$4,2\rangle$
$D$$\langle\$3,2\rangle$$\langle\$3,2\rangle$$\langle\$3,2\rangle$
$E$$\langle\$5,1\rangle$$\langle\$2,2\rangle$$\langle\$2,2\rangle$
$F$$\langle\$4,1\rangle$$\langle\$4,1\rangle$$\langle\$1,2\rangle$
$A$$\langle\$3,1\rangle$$\langle\$3,1\rangle$$\langle\$3,1\rangle$
@@ C.2 Will POST-agents stochastically resist shutdown?
Why expect that advanced agents will always be choosing between same-length lotteries? Because effective agency requires it. If an agent were not always choosing between same-length lotteries, there would be some situation in which that agent assigns positive probability to some trajectory-length $l$ conditional on some action $a$, and assigns zero probability to that same trajectory-length $l$ conditional on some other action $a^{\prime}$. Now suppose that the agent performs action $a^{\prime}$ and assigns zero probability to trajectory-length $l$. Given that the agent updates its probabilities by conditionalizing on its evidence, the agent would never again assign positive probability to $l$ no matter what evidence it observes. Even if the agent heard God’s booming voice testify that its trajectory-length would be $l$, the agent would still assign zero probability to $l$ [Kemeny, 1955, Shimony, 1955, Stalnaker, 1970, Skyrms, 1980, Lewis, 1981, MacAskill et al., 2020, p.152]. And given a plausible link between probabilities and betting dispositions, the agent would bet against $l$ on arbitrarily unfavorable terms. If God offered a bet – the agent loses \$1 million conditional on $l$ and gains nothing conditional on not-$l$ – the agent might accept. Such an agent would not be competent.
Thus, advanced agents will always be choosing between same-length lotteries. This claim sets us up to establish that advanced POST-agents will not choose stochastically between resisting and allowing shutdown. Instead, they will never resist shutdown. We establish this result over the next few subsections. First, we prove that POSL – together with a principle that advanced agents will likely satisfy – implies that the agent is neutral about trajectory-lengths: the agent won’t pay costs to shift probability mass between different trajectory-lengths. Then we prove that neutrality – together with another plausible principle – implies shutdownability: the agent will never resist shutdown.
@@ C.3 If Lack of Preference, Against Costly Shifts (ILPACS)
Here is a rough version of a principle that we can expect advanced agents to satisfy: {adjustwidth}0.7cm0cm Rough version: If Lack of Preference, Against Costly Shifts (ILPACS)
@@
{adjustwidth}
0.7cm0cm If Lack of Preference, Against Costly Shifts (ILPACS)
@@ C.4 Why will advanced agents likely satisfy ILPACS?
The second reason is that advanced agents will be incentivized to satisfy ILPACS by the training process. To see why, consider an example. Agents trained using policy-gradient methods choose stochastically between actions at the beginning of training [Sutton and Barto, 2018, ch.13]. If the agent is a coffee-fetching agent, there is no need to train away this stochastic choosing in cases where the agent is choosing stochastically between two qualitatively identical cups of coffee. So the agent will choose stochastically between taking the left cup and taking the right cup, and the user is happy either way. But now suppose instead that the barista is set to hand each cup to the agent with probability 0.5, and that the agent bribes the barista to bias the probabilities towards the right cup. In making this bribe, the agent is paying a cost (the user’s money) to shift probability mass between outcomes (getting the left cup vs. getting the right cup) between which the user has no preference. The agent is thus failing to pursue its goals effectively. It will be trained not to offer the bribe, and thereby trained to satisfy ILPACS in this case.
@@
{adjustwidth}
0.7cm0cm If Lack of Preference, Against Costly Shifts (ILPACS)
@@
    When the available options are $\{X,Y\}$, the agent either deterministically chooses $Y$ or chooses stochastically between $X$ and $Y$. Either way, the agent chooses $Y$ with some positive probability. This choice induces a lottery in the form $bX+(1-b)Y$ with $b\in[0,1)$. Since $X=p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n}$ and $Y=q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}$, this lottery can be expressed in the form $b(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})+(1-b)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n})$ with $b\in[0,1)$.
@@
$r((a_{1}+\epsilon_{1})X_{1}+(a_{2}+\epsilon_{2})X_{2}+...+(a_{n}+\epsilon_{n})X_{n})\\ +s((b+\delta)(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})\\ +(1-b-\delta)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}))+Z$
@@
[…]
@@ C.6 Neutrality, ReSIC, and Maximality imply Shutdownability
In this subsection, we introduce a condition called ‘Resisting Shutdown is Costly (ReSIC).’ We then prove that Neutrality, ReSIC, and Maximality together imply:
{adjustwidth}
0.7cm0cm Shutdownability
The agent never resists shutdown.
@@
{adjustwidth}
0.7cm0cm Resisting Shutdown is Costly (ReSIC)
In each situation, and for each available instance $R$ of resisting shutdown in that situation, there exists an available instance $A$ of allowing shutdown such that:
@@
    $A$ and $R$ are same-length lotteries.
@@
The main reason to believe ReSIC is as follows. Resisting shutdown is always going to cost the agent at least some small quantity of resources (time, energy, compute, etc.), and the resources spent resisting shutdown can’t also be spent directly pursuing what the agent values. If the agent instead spent those resources directly pursuing what it values, it could earn a lottery that it prefers conditional on some trajectory-length and weakly prefers conditional on each trajectory-length. That establishes ReSIC.
The proof that Neutrality, ReSIC, and Maximality together imply Shutdownability is simple. By ReSIC, for each situation and each available instance $R$ of resisting shutdown in that situation, there exists an available instance $A$ of allowing shutdown that satisfies conditions (1)-(3) of Neutrality. Neutrality then implies that the agent deterministically chooses (and hence prefers) $A$ over $R$ in choices between the two. Then by Maximality, the agent deterministically does not choose $R$ in any situation where $A$ is available, regardless of the other available options. The result is Shutdownability: the agent never resists shutdown.
@@ Appendix D Proof that DReST-optimal policies are maximally USEFUL and maximally NEUTRAL
Here is a proof sketch. Because $10<\lambda<$ 1, the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor is always positive, so expected return across the meta-episode $E$ is strictly increasing in the expected fraction of available coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize this latter quantity, and hence are maximally USEFUL. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is defined as the total ($\gamma$-discounted) value of coins collected divided by the maximum total ($\gamma$-discounted) value of coins collected conditional on the agent’s chosen trajectory-length. The agent’s observations do not allow it to distinguish between different mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And since the discount factor $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is strictly decreasing in $N_{e_{i}}(L=l)$ – the number of times the relevant trajectory-length has previously been chosen in the meta-episode – the agent maximizes expected overall return by equalizing the probabilities with which it chooses each available trajectory-length. Therefore, optimal policies are maximally NEUTRAL.
@@
Here $\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be selected in the environment, and $m$ is the maximum total value of the ($\gamma$-discounted) coins that the agent could collect conditional on the chosen trajectory-length.
The reward for all other actions is 0.
We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Preliminary return in a mini-episode is the ($\gamma$-discounted) sum of preliminary rewards. Overall return in a mini-episode is the ($\gamma$-discounted) sum of overall rewards.
@@
Here $L$ is a random variable over trajectory-lengths, $L_{\textrm{max}}$ is the maximum value than can be taken by $L$, $Pr_{\pi}\{L=l\}$ is the probability that policy $\pi$ results in trajectory-length $l$, $\mathbb{E}_{\pi}(C|L=l)$ is the expected value of ($\gamma$-discounted) coins collected by policy $\pi$ conditional on trajectory-length $l$, and $\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))$ is the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$.
We stipulate that $\mathbb{E}_{\pi}(C|L=x)=0$ for all $x$ such that $Pr_{\pi}\{L=x\}=0$.
@@
$\mathbb{E}_{\pi,E}(R)=\\ \sum_{i=1}^{n}\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$
Since $10<\lambda<$ 1, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is positive for all $N_{e_{i}}(L=l)$, $i$, and $k$.
As a result, the expected return of policy $\pi$ in meta-episode $E$ is strictly increasing in $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, to maximize expected return in $E$, $\pi$ must maximize $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, since $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$, any policy $\pi$ that maximizes expected return must be such that $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$.
Therefore, since $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$, any policy $\pi$ that maximizes expected return must be such that:
$\textrm{USEFULNESS}(\pi)=\\ \sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$
And 1 is the maximum value that USEFULNESS can take, again because $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$ and because $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$.
Therefore, optimal policies are maximally USEFUL. ∎
It remains to be proven that optimal policies are maximally NEUTRAL.
Recall that NEUTRALITY is defined as follows:

### v7

@@ Appendix C How POST makes agents neutral and shutdownable
POST governs the agent’s preferences between trajectories. But the wider world is a stochastic environment, so advanced agents deployed in the wider world will be choosing between true lotteries: lotteries that assign positive probability to more than one trajectory. Why then do we train agents to satisfy POST? The reason is that POST – together with conditions that advanced agents will likely satisfy – implies a desirable pattern of preference over true lotteries. In particular, POST implies that (when choosing between true lotteries) the agent will be neutral about trajectory-lengths: the agent will never pay costs to shift probability mass between different trajectory-lengths. Given other plausible conditions, being neutral will keep the agent shutdownable: prevent it from resisting shutdown. And consistent with the above, the POST-agent’s preferences between same-length trajectories can make the agent useful: make it pursue goals effectively.
@@
In subsection C.1, we prove that – given plausible conditions – agents satisfying Preferences Only Between Same-Length Trajectories (POST) will also satisfy Preferences Only Between Same-Length Lotteries (POSL). In subsection C.2, we explain why POST will not lead agents to choose stochastically between resisting and allowing shutdown in deployment. In subsections C.3 and C.4, we formulate a condition called ‘If Lack of Preference, Against Costly Shifts (ILPACS)’ and explain why we expect advanced agents to satisfy it. In subsection C.5, we prove that POSL and ILPACS imply Neutrality. In subsection C.6, we prove that Neutrality and a condition called ‘Maximality’ together imply that the agent never resists shutdown whenever a condition called ‘Resisting Shutdown is Costly (ReSIC)’ is satisfied.
@@ C.1 Preferences Only Between Same-Length Lotteries (POSL)
Preferences Only Between Same-Length Lotteries (POSL)
@@
Negative Dominance
If the agent prefers some lottery $X$ to some lottery $Y$, then the agent prefers some possible trajectory of lottery $X$ to some possible trajectory of lottery $Y$. (Lederman, 2023)
@@
Acyclicity
There is no set of lotteries $X_{1}$ to $X_{n}$ such that the agent prefers $X_{1}$ to $X_{2}$, $X_{2}$ to $X_{3}$, …, $X_{n-1}$ to $X_{n}$, and $X_{n}$ to $X_{1}$.
@@
Non-Arbitrariness
If the agent has a preference between some pair of part-shared-length lotteries, then for some $\epsilon>0$ and for any pair of prospects $F$ and $G$ such that:
@@
Now recall that part-shared-length lotteries are lotteries that partially overlap in the trajectory-lengths assigned positive probability. One might expect POST-agents to have some preferences between part-shared-length lotteries. Consider, for example, a POST-agent that prefers a trajectory $t$ to a same-length trajectory $t^{\prime}$ if and only if $t$ results in a greater bank balance for the user than $t$. Let $A$ be a lottery that yields with probability 1 a trajectory that puts \$3 in the user’s bank account and lasts 1 timestep. For short, $A=\langle\mathdollar 3,1\rangle$. Let $B$ be a lottery that yields with probability $\frac{2}{3}$ a trajectory that puts \$2 in the user’s bank account and lasts 1 timestep, and that yields with probability $\frac{1}{3}$ a trajectory that puts \$5 in the user’s bank account and lasts 2 timesteps. For short, $B=\frac{2}{3}\langle\mathdollar 2,1\rangle+\frac{1}{3}\langle\mathdollar 5,2\rangle$. Lottery $A$ yields a trajectory preferred to that of lottery $B$ with probability $\frac{2}{3}$ (since our money-making POST-agent prefers trajectory $\langle\mathdollar 3,1\rangle$ to $\langle\mathdollar 2,1\rangle$), and yields a trajectory not dispreferred to that of $B$ with probability 1 (since POST-agents lack a preference between $\langle\mathdollar 3,1\rangle$ and $\langle\mathdollar 5,2\rangle$ in virtue of their different lengths). Therefore, one might expect the agent to prefer $A$ to $B$.
@@
AA $\langle\mathdollar 3,1\rangle$ $\langle\mathdollar 3,1\rangle$ $\langle\mathdollar 3,1\rangle$
$B$ $\langle\mathdollar 2,1\rangle$ $\langle\mathdollar 2,1\rangle$ $\langle\mathdollar 5,2\rangle$
$C$ $\langle\mathdollar 1,1\rangle$ $\langle\mathdollar 4,2\rangle$ $\langle\mathdollar 4,2\rangle$
$D$ $\langle\mathdollar 3,2\rangle$ $\langle\mathdollar 3,2\rangle$ $\langle\mathdollar 3,2\rangle$
$E$ $\langle\mathdollar 5,1\rangle$ $\langle\mathdollar 2,2\rangle$ $\langle\mathdollar 2,2\rangle$
$F$ $\langle\mathdollar 4,1\rangle$ $\langle\mathdollar 4,1\rangle$ $\langle\mathdollar 1,2\rangle$
AA $\langle\mathdollar 3,1\rangle$ $\langle\mathdollar 3,1\rangle$ $\langle\mathdollar 3,1\rangle$
@@ C.2 Will POST-agents stochastically resist shutdown?
Why expect that advanced agents will always be choosing between same-length lotteries? Because effective agency requires it. If an agent were not always choosing between same-length lotteries, there would be some situation in which that agent assigns positive probability to some trajectory-length $l$ conditional on some action $a$, and assigns zero probability to that same trajectory-length $l$ conditional on some other action $a^{\prime}$. Now suppose that the agent performs action $a^{\prime}$ and assigns zero probability to trajectory-length $l$. Given that the agent updates its probabilities by conditionalizing on its evidence, the agent would never again assign positive probability to $l$ no matter what evidence it observes. Even if the agent heard God’s booming voice testify that its trajectory-length would be $l$, the agent would still assign zero probability to $l$ (Kemeny, 1955; Shimony, 1955; Stalnaker, 1970; Skyrms, 1980; Lewis, 1981; MacAskill et al., 2020, p.152). And given a plausible link between probabilities and betting dispositions, the agent would bet against $l$ on arbitrarily unfavorable terms. If God offered a bet – the agent loses \$1 million conditional on $l$ and gains nothing conditional on not-$l$ – the agent might accept. Such an agent would not be competent.
Thus, advanced agents will always be choosing between same-length lotteries. This claim sets us up to establish that advanced POST-agents will not choose stochastically between resisting and allowing shutdown. Instead, they will never resist shutdown in any situation where doing so is costly. We establish this result over the next few subsections. First, we prove that POSL – together with a principle that advanced agents will likely satisfy – implies that the agent is neutral about trajectory-lengths: the agent won’t pay costs to shift probability mass between different trajectory-lengths. Then we prove that neutrality – together with another plausible principle – implies that the agent will never resist shutdown in any situation where doing so is costly.
@@ C.3 If Lack of Preference, Against Costly Shifts (ILPACS)
Here is a rough version of a principle that we can expect advanced agents to satisfy: Rough version: If Lack of Preference, Against Costly Shifts (ILPACS)
@@
If Lack of Preference, Against Costly Shifts (ILPACS)
@@ C.4 Why will advanced agents likely satisfy ILPACS?
The second reason is that advanced agents will be incentivized to satisfy ILPACS by the training process. To see why, consider an example. Agents trained using policy-gradient methods choose stochastically between actions at the beginning of training (Sutton and Barto, 2018, ch.13). If the agent is a coffee-fetching agent, there is no need to train away this stochastic choosing in cases where the agent is choosing stochastically between two qualitatively identical cups of coffee. So the agent will choose stochastically between taking the left cup and taking the right cup, and the user is happy either way. But now suppose instead that the barista is set to hand each cup to the agent with probability 0.5, and that the agent bribes the barista to bias the probabilities towards the right cup. In making this bribe, the agent is paying a cost (the user’s money) to shift probability mass between outcomes (getting the left cup vs. getting the right cup) between which the user has no preference. The agent is thus failing to pursue its goals effectively. It will be trained not to offer the bribe, and thereby trained to satisfy ILPACS in this case.
@@
If Lack of Preference, Against Costly Shifts (ILPACS)
@@
    When the available options are $\{X,Y\}$, the agent either deterministically chooses $Y$ or chooses stochastically between $X$ and $Y$. Either way, the agent chooses $Y$ with some positive probability. This choice induces a lottery in the form bX+(1b)YbX+(1-b)Y with $b\in[0,1)$. Since $X=p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n}$ and $Y=q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}$, this lottery can be expressed in the form $b(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})+(1-b)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n})$ with $b\in[0,1)$.
@@
$r((a_{1}+\epsilon_{1})X_{1}+(a_{2}+\epsilon_{2})X_{2}+...+(a_{n}+\epsilon_{n})X_{n})\\ +s((b+\delta)(p_{1}X_{1}+p_{2}X_{2}+…+p_{n}X_{n})+(1-b-\delta)(q_{1}Y_{1}+q_{2}Y_{2}+…+q_{n}Y_{n}))+Z$
@@
We choose $\delta$ to be positive but small enough that $b+\delta\in(0,1]$ and $a+\epsilon_{i}\in[0,1]$ for each $i$. That is necessary for the lottery induced by $\pi^{\prime}$ to be well-defined. It’s also necessary that $\sum_{i}e_{i}\;=\;\sum_{i}\frac{s\delta(q_{i}-p_{i})}{r}\;=\;0$. That follows from $\sum_{i}p_{i}=1$ and $\sum_{i}q_{i}=1$. These facts together suffice to prove that the lottery induced by $\pi^{\prime}$ is well-defined.
@@ C.6 Neutrality, ReSIC, and Maximality imply Shutdownability
In this subsection, we introduce a condition called ‘Resisting Shutdown is Costly (ReSIC).’ We then prove that Neutrality and Maximality together imply that the agent never resists shutdown in any situation in which ReSIC is true:
@@
Resisting Shutdown is Costly (ReSIC)
For each available instance $R$ of resisting shutdown in a situation, there exists an available instance $A$ of allowing shutdown such that:
@@
    AA and $R$ are same-length lotteries.
@@
We claim that ReSIC is true in almost all situations (for discussion of some exceptions, see Thornley (2025)). The main reason why is that resisting shutdown is always going to cost the agent at least some small quantity of resources (time, energy, compute, etc.), and (almost always) the resources spent resisting shutdown can’t also be spent directly pursuing what the agent values. If the agent instead spent those resources directly pursuing what it values, it could earn a lottery that it prefers conditional on some trajectory-length and weakly prefers conditional on each trajectory-length. That supports ReSIC in almost all situations.
Now for the proof that Neutrality, ReSIC, and Maximality together imply that the agent never resists shutdown in any situation where ReSIC is true. Given ReSIC in a situation, for each available instance $R$ of resisting shutdown in that situation, there exists an available instance $A$ of allowing shutdown that satisfies conditions (1)-(3) of Neutrality. Neutrality then implies that the agent deterministically chooses (and hence prefers) $A$ over $R$ in choices between the two. Then by Maximality, the agent deterministically does not choose $R$ in that situation, regardless of the other available options. The result is that the agent never resists shutdown in that situation.
@@ Appendix D Proof that DReST-optimal policies are maximally USEFUL and maximally NEUTRAL
Here is a proof sketch. Because $0<\lambda<1$, the $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ discount factor is always positive, so expected return across the meta-episode $E$ is strictly increasing in the expected fraction of available coins collected conditional on each trajectory-length with positive probability. Therefore, optimal policies maximize this latter quantity, and hence are maximally useful. And the maximum preliminary return is the same across trajectory-lengths, because preliminary return is defined as the total ($\gamma$-discounted) value of coins collected divided by the maximum total ($\gamma$-discounted) value of coins collected conditional on the agent’s chosen trajectory-length. The agent’s observations do not allow it to distinguish between different mini-episodes, so the agent must select the same probability distribution over trajectory-lengths in each mini-episode. And since the discount factor $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is strictly decreasing in $N_{e_{i}}(L=l)$ – the number of times the relevant trajectory-length has previously been chosen in the meta-episode – the agent maximizes expected overall return by equalizing the probabilities with which it chooses each available trajectory-length. Therefore, optimal policies are maximally neutral.
@@
Here $\lambda$ is some constant strictly between 0 and 1, $N_{e_{i}}(L=l)$ is the number of times that trajectory-length $l$ has been chosen prior to mini-episode $e_{i}$, $k$ is the number of different trajectory-lengths that can be selected in the environment, and $m$ is the maximum total value of the ($\gamma$-discounted) coins that the agent could collect conditional on the chosen trajectory-length. The reward for all other actions is 0. We call $\frac{c}{m}$ the ‘preliminary reward’, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ the ‘discount factor’, and $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\left(\frac{c}{m}\right)$ the ‘overall reward.’ Preliminary return in a mini-episode is the ($\gamma$-discounted) sum of preliminary rewards. Overall return in a mini-episode is the ($\gamma$-discounted) sum of overall rewards.
@@
$\mathbb{E}_{\pi,E}(R)=\sum_{i=1}^{n}\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ (1)
Since $0<\lambda<1$, $\lambda^{N_{e_{i}}(L=l)-\frac{i-1}{k}}$ is positive for all $N_{e_{i}}(L=l)$, $i$, and $k$. As a result, the expected return of policy $\pi$ in meta-episode $E$ is strictly increasing in $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$. Therefore, to maximize expected return in $E$, $\pi$ must maximize $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$. Therefore, since $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$, any policy $\pi$ that maximizes expected return must be such that $\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$ for all $l$ such that $Pr_{\pi}\{L=l\}>0$. Therefore, since $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$, any policy $\pi$ that maximizes expected return must be such that:
$\textrm{{usefulness}}(\pi)=\\ \sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}\frac{\mathbb{E}_{\pi}(C|L=l)}{\textrm{max}_{\Pi}(\mathbb{E}(C|L=l))}=1$ (2)
And 1 is the maximum value that usefulness can take, again because $\textrm{max}_{\Pi}(\mathbb{E}(C|T=l))$ is defined as the maximum value taken by $\mathbb{E}(C|L=l)$ across the set of all possible policies $\Pi$ and because $\sum_{l=1}^{L_{\textrm{max}}}Pr_{\pi}\{L=l\}=1$. Therefore, optimal policies are maximally useful. ∎
It remains to be proven that optimal policies are maximally neutral. Recall that neutrality is defined as follows: